### v1

# Beyond Monolingual Assumptions: A Survey of Code-Switched NLP in the Era of Large Language Models

Abstract
Code-switching (CSW), the alternation of languages and scripts within a single utterance, remains a fundamental challenge for multiling ual NLP, even amidst the rapid advances of large language models (LLMs). Most LLMs still struggle with mixed-language inputs, limited CSW datasets, and evaluation biases, hindering deployment in multilingual societies. This survey provides the first comprehensive analysis of CSW-aware LLM research, reviewing ???? studies spanning five research areas, 12 NLP tasks, 30+ datasets, and 80+ languages. We classify recent advances by architecture, training strategy, and evaluation methodology, outlining how LLMs have reshaped CSW modeling and what challenges persist. The paper concludes with a roadmap emphasizing the need for inclusive datasets, fair evaluation, and linguistically grounded models to achieve truly multilingual intelligence. A curated collection of all resources is maintained at  [^1] https://github.com/lingo-iitgn/awesome-code-mixing/.
Beyond Monolingual Assumptions: A Survey of Code-Switched NLP in the Era of Large Language Models
[…]
1 Introduction
Code-switching (CSW), the alternation between two or more languages within a single utterance or discourse is a pervasive feature of multilingual communication worldwide (Poplack, 1988). With the rise of digital platforms, code-switched text has become ubiquitous across social media and online communication (Molina et al., 2016; Singh and Solorio, 2017), challenging NLP systems built on monolingual assumptions. In India, 26% of the population is bilingual and 7% trilingual (Chandramouli, 2011), with over 250 million speakers engaging in code-switched discourse, yet ASR systems trained on monolingual speech exhibit 30–50% higher word error rates on CSW data (Singh et al., 2025). Multilingual NLU models also suffer up to a 15% drop in semantic accuracy (Winata et al., 2021), leaving bilingual users underserved (Cihan et al., 2022). Figure 1 depicts intra- and inter-sentential code-mixing across multiple language pairs, emphasizing the linguistic variability that NLP systems must navigate.
Figure 1: Common model failures on code-mixed text: Takeaway(a) hallucination in MT translation (Bn-Hi-En), (b) factual inconsistency in POS tagging (Sp-En), and (c) misinterpretation in SA (Kz-Ru).
The evolution of CSW research mirrors key milestones in NLP. The Early Statistical Era (pre-2010) relied on rule-based and probabilistic models like n-grams, HMMs, and CRFs, laying the groundwork for bilingual text processing (Solorio and Liu, 2008). The Representation Learning Era (2010–2017) introduced distributed embeddings (Word2Vec) and recurrent models, advancing CSW tasks like LID, POS tagging, and NER (Solorio et al., 2014; Sequiera et al., 2015; Molina et al., 2016). The Contextual Understanding Era (2017–2020) brought GPT, BERT, XLM, and T5, enabling fine-tuning for code-switched data, though multilingual pretraining alone proved insufficient for robust CSW performance (Winata et al., 2021). The Foundation Model Era (2020–present) leverages LLMs like GPT-3, PaLM, and LLaMA for multilingual pretraining and prompt-based adaptation (Wang et al., 2025b).
LLMs have transformed CSW investigation across typologically diverse language pairs—Spanish-English (Sheokand et al., 2025), Hindi-English (Sheth et al., 2025), Chinese-English (Kong and Macken, 2025), Korean-English (Yoo et al., 2025), Ukrainian-Russian (Shynkarov et al., 2025), Cantonese-Mandarin (Dai et al., 2025), and Arabic-English (Issa et al., 2025), deepening our linguistic and sociocultural understanding of switching patterns (Yoo et al., 2024; Jehan et al., 2025). Methodologically, current LLMs support in-context mixing (Shankar et al., 2024), instruction tuning for low-resource CSW (Lee et al., 2024), synthetic data augmentation (Zeng, 2024), and advanced metrics evaluating structural and socio-pragmatic aspects (Ugan et al., 2025; Sterner and Teufel, 2025b).
Research Gap
Despite advances, LLMs struggle with zero-shot transfer to real-world CSW scenarios (Winata et al., 2023a). Multilingual LLMs often underperform compared to fine-tuned smaller models, showing that “multilingualism” alone does not ensure CSW proficiency (Zhang et al., 2023). LLMs also exhibit asymmetric performance: non-English tokens in English contexts degrade performance, while English tokens in other languages often enhance it (Mohamed et al., 2025), exacerbated by limited pretraining data for low-resource languages (Yoo et al., 2025).
This paper surveys the evolution of CSW research in the LLM era and presents a robust taxonomy (shown in Figure 2) that categorizes it into 5 key categories. The survey traces a transition from pre-LLM statistical and rule-based approaches focused on NLU tasks to LLM-era progress in NLG, speech, and multimodal domains, which show improved fluency but still struggle with factual consistency and low-resource performance. As summarized in Appendix §Table 1, specialized models, instruction tuning, and scalable architectures show notable gains, yet major gaps remain in low-resource languages, unseen code-mix patterns, and generation consistency. Future advances require diverse, semi-annotated datasets and CSW-aware evaluation metrics to better capture linguistic blending and contextual coherence.
Contributions
The key contributions include:
  • We provide the first comprehensive survey of CSW research in the LLM era, covering NLU and NLG tasks, datasets & benchmarks across diverse language pairs, real-world applications, and key architectural innovations in CSW-aware language processing.
  • This study reviews approximately ???? papers across five major CSW research areas, covering 12 core NLP tasks, analyzing over 30 datasets and benchmarks spanning more than 80 languages, and presenting a taxonomy (Figure 2) that classifies LLMs by architecture, training paradigm, and evaluations.
  • We identify key gaps in LLM-based CSW research, including limited low-resource support, script diversity challenges, weak cross-lingual transfer, and the absence of unified evaluation frameworks. Future progress depends on inclusive datasets, equitable models, and fair metrics.
2 Pre-LLM-Era works
Early computational approaches to code-switched word processing relied on rule-based and statistical models for foundational tasks such as language identification (LID) (Molina et al., 2016; Gundapu and Mamidi, 2018; Shekhar et al., 2020; Solorio and Liu, 2008; Chittaranjan et al., 2014; King et al., 2014), part-of-speech (POS) tagging (Vyas et al., 2014; Raha et al., 2019; Pratapa et al., 2018b; Sequiera et al., 2015), named entity recognition (NER) (Ansari et al., 2018; Singh et al., 2018b, a), and sentiment analysis (SA) (Patwa et al., 2020; Joshi et al., 2016). Methods included CRFs and CNNs for LID (Solorio and Liu, 2008; Chittaranjan et al., 2014), CNNs with n-grams for POS tagging (Vyas et al., 2014), character-level RNNs and SVMs for NER (Singh et al., 2018b), and SVM-based sentiment classification (Joshi et al., 2016). BiLSTM-CRF models with embeddings later improved LID and NER, reducing perplexity (Chopra et al., 2021; Zhang et al., 2023), while switch-point sampling enhanced Hinglish LID performance (Chatterjere et al., 2020). However, these approaches were limited by task and language-specific designs, shallow features, scarce labeled data, and poor cross-linguistic transfer (Molina et al., 2016; Shekhar et al., 2020). The fragmented nature of research led to isolated solutions, preventing the use of shared representations or unified frameworks across diverse CSW contexts (Winata et al., 2023a; Liu et al., 2022; Chi, 2025). The rise of LLMs has shifted CSW research toward unified, multilingual frameworks, motivating surveys to examine their adaptation, emerging trends, and future directions.
Code-Mixed Language AnalyticsTasksDatasets & ResourcesModel Training & AdaptationEvaluation & BenchmarkingMulti- & Cross-Modal ApplicationsNatural LanguageUnderstanding TasksNatural LanguageGeneration TasksDatasetsFrameworks& ToolkitsFine-tuning ApproachesPre-trainingApproachesPost-trainingApproachesBenchmarksEvaluation MetricsVision-LanguageProcessingSpeech ProcessingCross-ModalIntegration Sentiment & Emotion Analysis: Hinglish SentiMix (Baroi et al., 2020),Spanglish sentiment analysis (Patwa et al., 2020), EmoMix-3L (Bangla-English-Hindi) (Raihan et al., 2023b), multi-label emotion (Ameer et al., 2022; Dowlagar and Mamidi, 2021b) Syntactic Analysis: SyMCoM metrics (Kodali et al., 2022b), dependency parsing (Ghosh et al., 2019; Özateş et al., 2022) Named Entity Recognition: Cross-lingual NER (Zhou et al., 2022; Malmasi et al., 2022b; S and Shrivastava, 2022), entity boundary detection (Kannada-English pair)(Nandigam et al., 2022),(Hindi-English pair)(Singh et al., 2018a) POS Tagging: XLM-RoBERTa fine-tuning & transformer POS tagging(Absar, 2025),joint learning(Kartik et al., 2024; Chatterjee et al., 2022; Dowlagar and Mamidi, 2021c) Language Identification: word-level LID (Lambebo Tonja et al., 2022), Boundary detection (Sterner and Teufel, 2023), MaskLID (Kargaran et al., 2024), ECT (Kuwanto et al., 2024), LCT (Iliescu et al., 2021), Hope speech detection Hossain et al. (2021); Dave et al. (2021), Offensive language detection (Chakravarthi et al., 2021; Goswami et al., 2023; Jayanthi and Gupta, 2021; Balouchzahi et al., 2021a) Natural Language Inference: Cross-lingual NLI tasks (Chakravarthy et al., 2020; Shankar et al., 2024; Khanuja et al., 2020a; Prasad et al., 2021) Question Answering: MLQA(Lewis et al., 2020; Kumar et al., 2022b), XQuAD evaluation frameworks (Wu et al., 2025b), unified CSQA benchmark (Gupta et al., 2018; Chandu et al., 2018a) Intent Classification: Multilingual intent prediction and slot filling (Krishnan et al., 2021; Srivastava and Singh, 2020; Raihan et al., 2023a) Cross-lingual Transfer: Zero-shot transfer learning (Wu et al., 2023; Do et al., 2024; Lai et al., 2021; Whitehouse et al., 2022) Machine Translation: CS-to-English MT (Dhar et al., 2022; Kang, 2024; Borisov et al., 2025), pivot translation strategies (Huzaifah et al., 2024) Code-Mixed Text Generation: Synthetic data creation (Potter and Yuan, 2024; Salaam et al., 2022; Liu et al., 2022), LLM prompting (Kuwanto et al., 2024; Pratapa et al., 2018a) Text Summarization: Abstractive and extractive approaches (Mehnaz et al., 2021; Arora et al., 2023; Yoo et al., 2025) Transliteration: Indic language transliteration (Anand and Kumar, 2022; Vavre et al., 2022; Dowlagar and Mamidi, 2021b), script normalization (Nag et al., 2024) Dialogue Generation: Conversational agents(Hindi-English pair) (Agarwal et al., 2023; Liu et al., 2022; Kumar and Chakraborty, 2024) Multilingual Coverage: SwitchLingua (Xie et al., 2025),Megaverse (Ahuja et al., 2024),GLUECoS (Khanuja et al., 2020b) Low-Resource Coverage: Kazakh-Russian (KRCS) (Borisov et al., 2025; Balouchzahi et al., 2022; Goloburda et al., 2025),African (Ogunremi et al., 2023),Kanglish (Balouchzahi et al., 2022),Basque-Spanish (Heredia et al., 2025a),Moroccan Darija-French (Samih and Maier, 2016) Synthetic Data Generation: LLM-generated (Kuwanto et al., 2024), POS tagging Kartik et al. (2024), quality estimation (Hamed et al., 2025; Sravani and Mamidi, 2023; Srivastava and Singh, 2022a) Annotation Frameworks: COMMENTATOR (Sheth et al., 2024), ToxVidLM (Maity et al., 2024),Corpus (Kochar et al., 2024) Synthetic Data Generation Toolkits: Based on linguistic theories (Kuwanto et al., 2024; Gautam et al., 2021b), LLM-based generation (Potter and Yuan, 2024) Task-Adaptive Pre-Training (TAPT): Domain-specific adaptation and in-domain unlabeled data (Gururangan et al., 2020; Nayak and Joshi, 2022; Goswami et al., 2023) Specialized Code-Mixed Models: HingBERT (Nayak and Joshi, 2022), CoMix (Arora et al., 2023),COMMIT (Lee et al., 2024),SetFit Pannerselvam et al. (2024) Cross-lingual Alignment: COMMIT approach (Lee et al., 2024), continual learning (Zhang et al., 2021), embedding alignment (Lample et al., 2018) Instruction Tuning: (Lee et al., 2024; Wang et al., 2024; Zhang et al., 2025b) Multi-task fine-tuning: Joint training across CS tasks (Aguilar et al., 2020; Zhang et al., 2025b) Task-specific fine-tuning: Progressive complexity (Agro et al., 2025), switching point focus, easier-to-harder training (Yoo et al., 2025) Parameter Efficient Methods: QLoRA (Srivastava, 2025), knowledge distillation (Sidibomma et al., 2025) Reinforcement Learning for CSW Adaptation: RL with reward models & policy gradient methods (Palivela et al., 2025; Zhang et al., 2025c) Zero-, One- and Few-shot Learning: robustness improvement (Winata et al., 2021; Zhang et al., 2023),in-context learning Shankar et al. (2024); Held et al. (2023) Instance-Based Prompting: PRO-CS technique (Bansal et al., 2022; Udawatta et al., 2024; Deroy and Maity, 2025) Comprehensive Benchmarks: GLUECoS (Khanuja et al., 2020b), LinCE (Aguilar et al., 2020), MEGAVERSE (22 languages) (Ahuja et al., 2024), DravidianCodeMix (Chakravarthi et al., 2022) Domain-Specific Corpora: Social media (Chakravarthi et al., 2022), agricultural (Olaleye et al., 2025), conversational (Budzianowski et al., 2018) Task-Specific Metrics: chrF for MT (Popović, 2015), WER for ASR (Kadaoui et al., 2024), human evaluation protocols (Shimorina and Belz, 2022) CS-Specific Metrics: CMI (Chatterjere et al., 2020), SyMCoM (Kodali et al., 2022b) Traditional Metrics: F1-score (Raihan et al., 2023b), BLEU (Jasim et al., 2020), accuracy measures (Sterner and Teufel, 2025b) Quality Assessment: PhoBLEU (Arora et al., 2023), semantic preservation scores (Reimers and Gurevych, 2020), naturalness evaluation (Liu et al., 2021) Intrinsic Evaluation: Language modeling perplexity (Chen and Goodman, 1996), grammaticality scoring (Warstadt et al., 2019) Audio-Visual Recognition: CI-AVSR Cantonese (Dai et al., 2022), multimodal speech processing Raj Khan et al. (2021), multimodal speech processing Speech Translation: Direct speech-to-text translation (P S V N et al., 2025), End speech to speech (Weller et al., 2022), ASR and MoE system (Zhang et al., 2025a) End-to-End ASR: Code-switched speech recognition (Kronis et al., 2024), monolingual model fine-tuning (Bansal et al., 2022) Document Processing: Multilingual OCR (Dereza et al., 2024), code-switched document analysis (Do et al., 2024) Multimodal VQA: Knowledge distillation approaches (Raj Khan et al., 2021), code-mixed visual question answering (Ahuja et al., 2024) Phonetic Processing: Transliteration modules (Weisberg Mitelman et al., 2024), phonetic embedding for ASR (Chi and Bell, 2022) Multimodal Fusion: Audio-visual-text systems for hate speech (Srivastava, 2025), cross-modal alignment for emotion (Wu et al., 2025a)
Figure 2: A taxonomy of the code-switching research landscape. Takeaway The mind map highlights important works across the diverse categories of code-switching research.
3 Popular CSW Tasks
3.1 Natural Language Understanding Tasks
Language Identification
Word & sentence-level LID form the foundation of CSW NLP. Transformer-based models substantially improve LID accuracy, with TongueSwitcher enhancing boundary detection in German–English CSW and outperforming traditional baselines like FastText and CLD (Lambebo Tonja et al., 2022; Sterner and Teufel, 2023). MaskLID identifies subdominant language patterns, while integrating Equivalence Constraint Theory with transformers enables detection of grammatically valid switch points (Kargaran et al., 2024; Kuwanto et al., 2024). Beyond pure LID, related efforts extend to hope and offensive speech detection—addressing multilingual and adversarial switching challenges—using datasets in Roman-Urdu and Malayalam-English (Ahmad et al., 2025; Balouchzahi et al., 2021b), as well as models like COOLI and SetFit are for efficient zero and few-shot approaches (Balouchzahi et al., 2021a; Pannerselvam et al., 2024).
Part-of-Speech Tagging
LLM-based approaches improve CSW POS tagging via pre-trained models, multi-task learning, and transfer strategies. mBERT embeddings capture Arabic–English CSW nuances, and character-CNNs with mBERT further boost Hinglish POS (Sabty et al., 2020; Aguilar and Solorio, 2020). Joint LID–POS transformer models and switch-point–biased self-training improve accuracy on Telugu–English data (Dowlagar and Mamidi, 2021c; Chopra et al., 2021). Bilingual pretraining on GLUECoS enhances POS (Winata et al., 2021; Prasad et al., 2021). PACMAN supports fine-tuning, while PRO-CS and CoMix leverages prompts for improved POS tagging and generation of code-switched text (Chatterjee et al., 2022; Kumar et al., 2022b; Bansal et al., 2022; Arora et al., 2023). XLM-R fine-tuned with the S-index achieves state-of-the-art on Hinglish POS improves prformance (Absar, 2025).
Named Entity Recognition
NER in CSW contexts has evolved significantly, from using embedding attention for Spanglish tweets (Wang et al., 2018), to multilingual frameworks like MELM that leverage synthetic CSW pretraining (Zhou et al., 2022). Two-stage CMB models improve Hinglish NER by separating span detection and classification (Pu et al., 2022). Benchmarks such as MultiCoNER and toolkits like CodemixedNLP support multilingual and Hinglish NER (Malmasi et al., 2022b; Jayanthi et al., 2021). Contextualized embeddings aid Arabic–English NER (Sabty et al., 2020), while pseudo-labeling and switch-point–biased self-training enhance boundary detection (El Mekki et al., 2022; Chopra et al., 2021). Prompt-based models like PRO-CS and data augmentation in CoSDA-ML enable zero-shot transfer (Bansal et al., 2022; Qin et al., 2020). Generative frameworks such as GPT-NER achieve strong few-shot performance but suffer from hallucinations (Wang et al., 2025a), whereas GLiNER outperforms ChatGPT in zero-shot settings (Zaratiana et al., 2024). Persistent challenges in NER include limited data, complex entities, and trilingual ambiguity (Dowlagar and Mamidi, 2022; Malmasi et al., 2022b).
Takeaway
Although notable progress has been made in core CSW NLU tasks, many frontier areas remain underexplored, highlighting opportunities to extend research beyond existing linguistic and computational paradigms.
3.2 Natural Language Generation Tasks
Code-Mixed Text Generation
Semi-supervised transfer learning and translation models improved Hinglish fluency (Gupta et al., 2020; Tarunesh et al., 2021), while COCOA achieved strong English–Spanish generation (Mondal et al., 2022). Dependency tree methods produced CSW text without parallel corpora (Gregorius and Okadome, 2022). Synthetic filtering and LLM prompting enhanced Tagalog–English generation (Sravani and Mamidi, 2023; Yong et al., 2023; Terblanche et al., 2024), and LLMs aided grammatical correction (Potter and Yuan, 2024; Heredia et al., 2025b). However, EZSwitch and HinglishEval revealed weak alignment between metrics and human judgments (Kuwanto et al., 2024; Srivastava and Singh, 2022a).
Machine Translation
Back-translation improved Hinglish MT, while unsupervised approaches enhanced Sinhala–English (Tarunesh et al., 2021; Kugathasan and Sumathipala, 2021). COMET-based synthetic data boosted Indic translation (Gupta et al., 2021a; Gautam et al., 2021b). Gated seq2seq with language tags and fine-tuned mT5/mBART advanced Hinglish and MSA–Egyptian MT (Dowlagar and Mamidi, 2021a; Jawahar et al., 2021a; Nagoudi et al., 2021). LLM syntactic post-processing improved Cantonese–Mandarin translation (Dai et al., 2025), and GPT-3 prompting enhanced Hinglish fluency (Khatri et al., 2023). Fine-tuned T5 achieved strong CodeMix-to-English results (Chatterjee et al., 2023; Jawahar et al., 2021b; Raj Khan et al., 2021). Persistent challenges in MT include syntactic misalignment and LLM limitations (Winata et al., 2021; Sazzed, 2021).
Takeaway
The LLM era has advanced NLG tasks by enabling generation and conversational capabilities that pre-LLM approaches could not achieve, though progress remains largely confined to high-resource pairs, with English-centric biases, limited low-resource corpora, typological diversity, and syntactic complexity still hindering generalization.
A detailed discussion of remaining NLU and NLG tasks is provided in Appendix §A.1 and §A.2.
4 Datasets and Resources
4.1 Datasets
Multilingual Coverage
Large-scale code-switching corpora, pre-trained on mixed-language text (e.g., millions of tokens from diverse sources), enhance NLU transfer via synthetic augmentation (Zhang et al., 2024a). The Multilingual Identification of English Code-Switching dataset, manually annotated token-level classifiers on high-resource pairs, benchmarks English switches across unseen languages (Sterner, 2024). SwitchLingua, a manually curated multi-ethnic collection, spans 420k textual samples and >80 hours audio across 12 languages/63 ethnic groups, reducing bias via LLM-assisted balancing (Xie et al., 2025). MEGAVERSE, an LLM-driven synthetic/manual benchmark, covers 22 datasets in 83 languages with translation/alignment for multimodal evaluation (Ahuja et al., 2024). MultiCoNER, manually annotated across 3 domains/11 languages (expanded to 12 in v2 with 33 entity classes), uses LLM synthetic augmentation for code-mixed NER (Malmasi et al., 2022b). NusaX, human-annotated parallel sentiment corpus, provides data for 10 Indonesian local languages to support indirect CSW analysis (Winata et al., 2023b). GLOSS, built on pre-trained MT models with a code-switching module, synthesizes CSW texts for absent language pairs, enabling generalization without manual curation (Hsu et al., 2023). This LLM-centric shift enables scalable, inclusive benchmarks, though typological gaps persist.
Low Resource Coverage
Low-resource datasets includes BnSentMix, a manually annotated Bengali-English dataset (20,000 samples with 4 sentiment labels) (Alam et al., 2025) fine-tuned on XLM-R, DravidianCodeMix - manually annotated for Tamil-English (SA/LID: 64,773/64,991 samples), Kannada-English (SA/LID: 30,872 / 29,601 samples), and Malayalam-English (SA/LID: 9,751 / 8,983 samples), supports SA and offensive LID (Chakravarthi et al., 2022), COMMIT uses LLM-based instruction tuning with synthetic CSW data, blocking English and using five low-resourced languages: Greek, Hindi, Thai, Tamil, and Bengali, aligning via causal language modeling, for QA (Lee et al., 2024), Marathi–English corpora manually annotated with pre-trained models, boosts NER and SA tasks(Joshi et al., 2023), SentMix (Bangla-English-Hindi trilingual for NLI) (Raihan et al., 2023a), and GPT-3.5 synthetic Afrikaans/Yoruba–English data (Terblanche et al., 2024) improves zero-shot sentiment accuracy. These efforts diversify CSW NLP by reducing dependence on high-resource language pairs.
Synthetic Data Generation and Augmentation
The scarcity of annotated code-mixed datasets has driven synthetic data generation to enhance multilingual NLP. A dependency parser for Bengali-English, augmented with a synthetic treebank (270,531 sentences), improved parsing by UAS (Winata et al., 2019). PhraseOut substitutes high-confidence phrases for multilingual NMT, on Hindi-English (Jasim et al., 2020). A semi-supervised approach with pre-trained encoders generated Hindi-English CSW text, (Gupta et al., 2020). CoSDA-ML’s synthetic multilingual data fine-tuned on mBERT, improves zero-shot NLI across 19 languages (Qin et al., 2020). mBERT-based augmentation with ternary sequence labeling enhanced Hindi-English MT (Gupta et al., 2021a). VACS, a variational autoencoder, generated synthetic CSW text achieves reduced perplexity (Samanta et al., 2019). LLM-based generation for Hindi-English puns and sentiment (49,560 synthetic corpus in Spanish-English & Malayalam-English) improved BLEU score (Zeng, 2024; Sarrof, 2025). In-Context Mixing (ICM) prompts enhanced intent classification (Shankar et al., 2024). SynCS datasets (15,500 parallel sentences) aligned representations, boosting zero-shot performance % (Wang et al., 2025b). A naturalistic CSW dataset with parallel translations improved PLM evaluation (Leon et al., 2024). However, human judgments reveal that synthetically generated CSW sentences often lack naturalness, with only 60-65% acceptability compared to human-authored text (Kodali et al., 2025).
Takeaway
LLM-generated datasets expand low-resource coverage but often lack naturalness and cultural nuance. Semi-automated, human-in-the-loop annotation can yield more authentic and scalable resources. (See Table 6 in the Appendix §A.5) for the list of datasets covered).
4.2 Frameworks and Toolkits
Annotation Frameworks
High-quality annotations are crucial for effective supervised models in code-mixed NLP. CoSSAT enables speech annotation (Shah et al., 2019), while COMMENTATOR integrates LLMs as an API for robust text annotation to support prediction alongside annotation. (Sheth et al., 2024). CHAI uses RL from AI feedback to enhance CSW translation annotations (Zhang et al., 2025c) and multimodal frameworks like ToxVidLM (Maity et al., 2024) manage complex video toxicity detection in CSW.
Synthetic Data Generation Toolkits
To address persistent data scarcity, toolkits such as GCM (Rizvi et al., 2021), utilized by Huzaifah et al. (2024) for generating valid CSW text, and CodemixedNLP, an open-source library designed for creating synthetic CSW texts across seven Hinglish tasks (Jayanthi et al., 2021) offer robust solutions for enhancing CSW text generation. These tools facilitate large-scale corpus creation for tasks such as MT (Sravani and Mamidi, 2023) and SA (Zeng, 2024).
Takeaway
Synthetic data generation for annotation workflows accelerates CSW research, but the scarcity of multilingual tools limits impact, highlighting the need for adaptable, low-resource–focused toolkits to build natural and useful benchmarks for LLM training and evaluation.
5 Model Training & Adaptation
5.1 Pre-training Approaches
Specialized Code-Mixed Models
HingBERT built on L3Cube-HingCorpus outperforms mBERT on GLUECoS (Nayak and Joshi, 2022), while code-mixed embeddings improve hate speech detection over monolingual ones (Pratapa et al., 2018b; Banerjee et al., 2020). HingBERT variants enhance NLP tasks in zero-shot settings (Shirke et al., 2025; Patil et al., 2023), and natural CSW data outperforms synthetic alternatives (Santy et al., 2021). For multimodal tasks, CMCLIP surpasses baselines by 5–10% in image-text retrieval (Kumari et al., 2024), and CMLFormer’s dual-decoder improves CSW modeling (Baral et al., 2025; Suresh et al., 2024), with hierarchical transformers enhancing long-sequence handling (Suresh et al., 2024).
Task-Adaptive Pre-Training
CoMix boosts translation by 12.98 BLEU (Arora et al., 2023), boundary-aware MLM improves QA/SA across Hindi-English, Spanish-English, Tamil-English, and Malayalam-English (Das et al., 2023), and alignment-based pre-training yields +7.32% SA, +0.76% NER, +1.9% QA for Hinglish (Fazili and Jyothi, 2022; Prasad et al., 2021). Multilingual augmentation enhances intent classification (Krishnan et al., 2021), and SynCS achieves +10.14 points in Chinese language while outperforming 20x equivalent monolingual data (Wang et al., 2025b).
Cross-lingual Alignment
CoSwitchMap improves bilingual lexicon induction (Gaschi et al., 2023), synthetic CSW achieves 5.1 MRR@10 for cross-lingual retrieval (Litschko et al., 2023), and CMLFormer captures language transitions better (Baral et al., 2025). MVMLT enhances NER alignment (Lai et al., 2021), AIMLT improves intent detection by 4–6% (Zhu et al., 2023; Micallef et al., 2024), and X-CIT fine-tunes Llama-2-7B across five languages (Wu et al., 2025b).
Takeaway
Pretraining approaches have improved CSW performance, with specialized architectures like CM-CLIP and HingBERT enabling better handling of multilingual and multimodal inputs, though synthetic data may still miss nuances.
5.2 Mainstream Fine-tuning Approaches
Task-specific fine-tuning
Fine-tuning transformers achieve state-of-the-art performance on Kannada-English LID (Lambebo Tonja et al., 2022), fine-tuned XLM-RoBERTa for POS tagging introduces the S-index for measuring switching intensity and demonstrating effective generalization (Absar, 2025). In SA and dialogue, transformer and multi-task models improve robustness on noisy social media data (Palomino and Ochoa-Luna, 2020; Wu et al., 2020). MT also benefits from fine-tuned mBART, mT5, or custom models, enhancing fluency for Hinglish and related languages (Chatterjee et al., 2023; Khan et al., 2022). For Cantonese–Mandarin speech translation, a three-component fine-tuning strategy improves BLEU scores (Dai et al., 2025), while Yoruba-English ASR fine-tuning with FastConformer Transducer reduces WER from 62.94% to 14.81% using 13× fewer parameters than Whisper Large v3 (Babatunde et al., 2025).
Multi-task fine-tuning
Intermediate-task mBERT fine-tuning improves NLI, QA, and SA on GLUECoS by 2–3% for Hinglish and Spanish-English (Prasad et al., 2021). Joint training boosts offensive LID on Hinglish tweets and enhances NER and code-switch detection in Algerian-Arabic (Prasad et al., 2021; Amazouz et al., 2017), multi-directional fine-tuning raises Hinglish-English translation BLEU by 40% (Kartik et al., 2024), and modular approaches like AdapterFusion, prompt tuning, and ContrastiveMix further improve zero-shot IR and transfer (Rathnayake et al., 2024; Do et al., 2024).
Takeaway
Fine-tuning LLMs on CSW corpora boosts NLP task performance via shared representations, curriculum learning, and synthetic data, but models still struggle with unseen switches, low-resource languages, syntactic variation, and spontaneous CSW.
A detailed discussion of remaining Fine-tuning approaches is provided in Appendix §A.5.
5.3 Post-training Approaches
Zero-, One- and Few-shot Learning
Zero-shot methods include prompting LLMs like GPT-3.5 for CSW generation in Southeast Asian languages (Yong et al., 2023), EntityCS’s entity-level CSW with Wikidata for cross-lingual slot filling (10% gain) (Chen et al., 2022), MulZDG’s framework for dialogue generation (Liu et al., 2022), MALM’s augmented pre-training for MT (Gupta, 2022), saliency-based multi-view mixed language training for classification (Lai et al., 2021), multilingual CSW augmentation for intent/slot filling (4.2% accuracy gain) (Krishnan et al., 2021), and CoSDA-ML’s data augmentation for cross-lingual NLP (0.70 score) (Qin et al., 2020); however, LLMs like GPT-4 show 14-point accuracy drops in zero-shot CSW tasks, with pretraining language choice impacting performance significantly in such settings (Zhang et al., 2023; Tatariya et al., 2023). One and Few-shot methods include adapted prompting with ChatGPT for low-resource understanding via semantic similarity-based demonstration selection (Tahery and Farzi, 2025), RAG-based in-context learning (1-8 examples) for hate speech detection (63.03 $\mu$-F1 with GPT-4o) (Srivastava, 2025), multi-task LLM fine-tuning for harmful content in memes (Indira Kumar et al., 2025), generative transformers for emotion detection in Bangla-English-Hindi (Goswami et al., 2023), and translation/LLM classification for affective tasks (Yadav et al., 2025); one-shot is typically grouped with few-shot, as in RAG-retrieved examples boosting F1 by 20 points (Srivastava, 2025). These methods rely on prompting, data augmentation, and entity-centric switching, but unseen language pairs and low-resource scenarios remain challenging.
Instance-based Prompting
Instance-based prompting in CSW tasks enhances LLM performance, with PRO-CS using mBERT with Hinglish prompts and improving NER and POS tagging by 10–15% (Bansal et al., 2022). GLOSS synthesizes CSW text for unseen pairs and achieving 55% BLEU/METEOR gains through self-training (Hsu et al., 2023). DweshVaani’s RAG retrieves 1–8 Hinglish examples, boosting hate speech detection to 63.03 score on Gemma-2 (Srivastava, 2025). In-Context Mixing improves intent classification by 5–8% on MultiATIS++ (Shankar et al., 2024). Despite these gains, informal CSW and low-resource language pairs remain challenging (Bansal et al., 2022; Shankar et al., 2024).
Takeaway
Post-training methods enable LLMs to refine their knowledge, improve reasoning, enhance factual accuracy, and align more effectively with user intents and ethical considerations in CSW context.
6 Evaluation & Benchmarking
6.1 Benchmarks
Domain-Specific & Comprehensive Benchmarks
CodeMixBench, with 5k+ Hinglish, Spanglish, and Chinese Pinyin–English prompts, shows 5–10% drops over English-only tasks using fine-tuned CodeLLaMA (Sheokand et al., 2025), while MEGAVERSE spans 22 datasets and 83 languages with LLM translation and LoRA adapters for low-resource QA (Ahuja et al., 2024). Domain-targeted corpora like the Telugu–English medical dialogue dataset (3k dialogs, 29k utterances) leverage XLM-R for intent and slot filling (Dowlagar and Mamidi, 2023), and MultiCoNER covers 11 languages with LLM augmentation and multi-task transformers, improving over mBERT (Malmasi et al., 2022b). SwitchLingua contributes 420k texts and 80+ hours of multilingual audio using LLM-assisted balancing and LoRA fine-tuning (Xie et al., 2025).
Comprehensive benchmarks such as GLUECoS (Khanuja et al., 2020b) and LinCE (Aguilar et al., 2020) enable multi-task LLM evaluation, while CroCoSum (Zhang and Eickhoff, 2024) and DravidianCodeMix (Chakravarthi et al., 2022) provide manually annotated data for summarization, SA, and offensive LID tasks. PACMAN (Chatterjee et al., 2022) and COMI-LINGUA (Sheth et al., 2025) reduce annotation effort through semi-automated methods. These benchmarks combine synthetic generation, semi-automated annotation, and efficient architectures like adapters for robust, scalable CSW evaluation.
Takeaway
Evaluation resources are diverse, and while synthetic augmentation can accelerate CSW research, future efforts should focus on multilinguality, multi-modality and generative tasks (See Table 5 in the Appendix A.5 for the list of benchmarks covered).
6.2 Evaluation Metrics
Traditional Metrics
Early CSW NLP evaluations used monolingual metrics like F1, Precision, Recall, Accuracy for classification (Qin et al., 2020; El Mekki et al., 2022), and BLEU, ROUGE, METEOR for generative tasks (Agarwal et al., 2021; Gupta et al., 2022), but often miss CSW complexities due to rigid lexical matching (Papineni et al., 2002; Lin, 2004; Hada et al., 2024).
CS-Specific Metrics
CSW-specific metrics address these limitations. Code-Mixing Index (CMI) measures word-level mixing (Das and Gambäck, 2013), SyMCoM evaluates grammatical well-formedness (Kodali et al., 2022a), I-Index assesses switch-point integration (Guzmán et al., 2017), and Switch Point-based Metrics analyze intra- and inter-sentential switches (Gambäck and Das, 2016). HinglishEval combines linguistic metrics with embeddings for better quality estimation (Kodali et al., 2022b; Srivastava and Singh, 2022a).
Task-Specific Metrics
Semantic-Aware Error Rate (SAER) enhances ASR evaluation (Xie et al., 2025), PhoBLEU handles orthographic variations (Arora et al., 2023), and chrF++ evaluates character n-grams for MT in morphologically rich languages (Popović, 2015). For Quality Assessment, Metrics like SyMCoM and PIER (Kodali et al., 2022a; Ugan et al., 2025) assess syntactic and semantic quality, while Cline emphasizes human-judged naturalness (Kodali et al., 2025).
Intrinsic Evaluation
Methods like distinguishing ground-truth from phonetically similar sentences (Chen and Goodman, 1996) and the gold-standard-agnostic GAME metric (Gupta et al., 2024) offer flexible CSW evaluation.
Takeaway
CSW evaluation has evolved from monolingual metrics to CS-specific metrics but more linguistically informed, code-mixing aware metrics are needed to assess fairness, safety, and cross-lingual robustness in multilingual models.
7 Multi- & Cross-Modal Applications
7.1 Speech Processing
Speech Translation & End-to-End ASR
CSW speech translation has progressed with end-to-end models for English–Spanish (Weller et al., 2022) and real-world streaming Mandarin–English via self-training (Alastruey et al., 2023). Speech segmentation models, such as Whisper fine-tuned on CoVoSwitch Korean–English CSW data (Kang, 2024), and linguistic-informed data augmentation have improved ASR performance (Chi and Bell, 2022; Hamed et al., 2025). Monolingual fine-tuning outperformed multilingual models for Yoruba–English (Babatunde et al., 2025), while Hindi-Marathi datasets enabled transformer ASR (Hemant and Narvekar, 2025), Wav2Vec2 and GPT-2 enhance transcription accuracy in multilingual indian contexts with CSW pairs and dialectal variations (R et al., 2025), and MoE with SC-LLMs further boosted Mandarin-English ASR in CSW (Zhang et al., 2025a).
Audio-visual Recognition
Enhanced Visual features in Mandarin-English and low-resource Yoruba-English ASR (Babatunde et al., 2025), while Hindi-Marathi CS transformers and Wav2Vec2 fusion aided Indian languages like Tamil (Hemant and Narvekar, 2025).
7.2 Vision-Language Processing
Multimodal VQA & Document Processing
Multimodal VQA for CSW advanced via knowledge distillation, enables Hinglish systems to handle CS queries (Raj Khan et al., 2021). LLMs detected harmful content in code-mixed memes (Indira Kumar et al., 2025), BanglAssist used RAG for Bengali-English CSW (Kruk et al., 2025), and Kazakh-Russian LLMs evaluated safety with region-specific prompts (Goloburda et al., 2025). LLM agents analyzed Sinhala-English and Tamil-English CSW interviews (Zhao et al., 2025).
Document processing employed multilingual OCR for robust extraction (Dereza et al., 2024) and contrastive learning improved Vietnamese-English analysis (Do et al., 2024).
7.3 Cross-Modal Integration
Phonetic & Multimodal Processing
Phonetic modeling enhances multilingual CSW performance through discriminative language modeling (Winata et al., 2019), transliteration for agglutinative languages (Tasawong et al., 2023), and applications such as abusive content detection (Gautam et al., 2021a), translation alignment (Chou et al., 2023), and back-transliteration (Fernando and Ranathunga, 2021). Transformer-based phonetic guidance (Yang and Tu, 2022) and Wav2Vec2–GPT-2 fusion (Perera and Sumanathilaka, 2025) further improve CSW phonetic tasks. Multimodal fusion integrates audio, visual, and textual cues, enhancing code-mixed ASR (Tamil, Kannada, Mandarin-English) and toxicity detection in videos (Maity et al., 2024; Perera and Sumanathilaka, 2025; Zhang et al., 2025a).
Takeaway
Multimodal, phonetic, and end-to-end approaches are advancing CSW speech, vision, and document processing, but robust integration across low-resource languages and modalities remains a key challenge.
8 Emerging Frontiers in CSW Research
Futuristic Directions in CSW NLP
Futuristic Datasets
The progress of CSW NLP relies on developing expansive and inclusive datasets but large-scale conversational resources capturing dynamic CSW interactions remain absent, and the lack of human-preference datasets limits adaptive modeling. Multimodal efforts such as MEGAVERSE (Ahuja et al., 2024) are promising but still restricted in linguistic and domain coverage. Building diverse corpora, such as Hindi–Marathi speech (Hemant and Narvekar, 2025) or a large multi-domain multilingual dialogue dataset (Moradshahi et al., 2023), faces persistent challenges of cost and scalability. Future work must integrate synthetic data generation with expert validation to achieve diversity and quality, particularly for underrepresented languages.
Futuristic Architectures
Next-generation architectures should integrate CSW text, speech, and vision to improve switch-point detection, contextual understanding, and natural multilingual interactions, while ASR and TTS systems leverage self-supervised encoders, cross-lingual, and emotion-aware conditioning.
Futuristic Evaluation
Futuristic CSW-aware evaluation will move beyond isolated metrics to holistic, human-aligned assessment of multilingual competence, jointly measuring switch-point prediction, semantic consistency, and fluency. Adaptive scoring models trained on human preferences will capture nuanced judgments of naturalness and cross-lingual balance.
Futuristic CSW-aware Ethical AI
It will prioritize inclusivity and transparency, using bias-aware training for equitable performance across languages and dialects. Fairness metrics, participatory data curation, and transparent documentation will ensure accountability and foster socially and linguistically fair CSW systems.
Futuristic Applications
CSW research can drive high-impact applications such as multilingual conversational assistants for accessible public services, cross-lingual education platforms that adapt to learners’ language mixes, and digital preservation tools for endangered dialects. These applications are socially and economically significant, reducing language barriers, enhancing equitable access to knowledge and services, and empowering communities in the multilingual digital economy.
9 Conclusion
CSW research has undergone a major transformation with the rise of LLMs. From early rule-based and statistical methods to multilingual pretraining, fine-tuning, and instruction-based adaptation, the field has steadily moved toward more unified and scalable approaches. This survey shows that while significant progress has been made in LID, tagging, translation, and generation for high-resource pairs, major gaps persist for low-resource and culturally diverse contexts. The future of CSW research lies in building inclusive datasets, improving evaluation protocols, and designing socially grounded systems that reflect the realities of multilingual communication.
Limitations
Despite providing a broad survey, this paper has several limitations:
  1. 1.
    Coverage Bias The survey highlights widely studied language pairs and might have missed indigenous or minority code-mixed languages.
  2. 2.
    Evolving Landscape Given the rapid pace of LLM research, some approaches and benchmarks described may soon be outdated or replaced by newer paradigms.
  3. 3.
    Evaluation Constraints The analysis is largely focused on text-based tasks, whereas speech, multimodal, and conversational CSW applications remain underexplored.
  4. 4.
    Practical deployment The survey mainly covers academic progress, leaving ethical, computational, and accessibility concerns in real-world deployment less examined.
Ethics Statement
This study involves a review and synthesis of previously published research and publicly available datasets. No human or user data were collected or analyzed. All works included in this survey were cited appropriately to acknowledge original authorship. The review process was conducted with transparency and fairness, avoiding selective reporting or biased interpretations. Our study promotes fairness and inclusivity in multilingual NLP by focusing on underrepresented code-mixed language scenarios, encouraging equitable research attention toward linguistically diverse communities. The study adheres to established ethical standards for research in computational linguistics.
Appendix A Methodology
This section outlines the methodology adopted to identify, review, and categorize literature relevant to this survey on code-switching in the era of LLMs. The goal was to capture key trends, modeling techniques, datasets, and challenges across NLP tasks rather than conduct an exhaustive systematic review. The approach follows established survey practices (Kinney et al., 2023).
Paper Selection
We began by defining a set of search keywords targeting three core dimensions: code-mixing/code-switching, multilingual NLP, and large language models. To ensure broad linguistic coverage, the search included major bilingual and multilingual language pairs documented in prior works and repositories (e.g., CoVoSwitch, GLUECoS, LinCE). Using these keywords, we queried the ACL Anthology, arXiv and Semantic Scholar databases via their APIs, with a search cutoff date of October 2025, consistent with ACL Rolling Review’s recency guidelines. This process initially retrieved  500 papers.
Screening and Filtering
Duplicate entries were removed using DOIs and titles, prioritizing peer-reviewed sources. The remaining papers were manually screened for relevance. A study was included if it addressed code-mixing or code-switching within any NLP task, or explored multilingual and LLM-based adaptation methods. This screening resulted in a refined set of 304 papers, covering both pre-LLM and LLM-era research.
Categorization
Selected papers were categorized by (i) task type (e.g., Language Identification, POS Tagging, NER, Sentiment Analysis, Speech/ASR, MT, NLG), (ii) modality (text, speech, vision-language), and (iii) model architecture (transformer-based, instruction-tuned, multimodal). When overlaps occurred (e.g., between translation and generation), we retained the category most central to the contribution. Dataset coverage, annotation methods, and language pairs were systematically verified to map diversity and resource availability. High-, mid-, and low-resource classifications followed conventions in multilingual NLP research.
This multi-stage process of search, screening, and categorization produced  ???? papers forming the foundation of this survey, spanning 12 NLP tasks, 30+ datasets, and 80+ languages.
A.1 Natural Language Understanding Tasks
Language Identification
Script detection remains crucial for accurate token-level processing, with Bi-GRU architectures achieving 90.17% accuracy on Roman Urdu, Hindi, Saraiki, Bengali, and English using GloVe embeddings (Yasir et al., 2021). The ILID corpus provides 250K sentences across 25 scripts and 23 languages, including dual-script instances for Manipuri and Sindhi (Ingle and Mishra, 2025). Character n-gram TF-IDF features (1–6 grams) have proven effective for Dravidian script-mixed social media text (Saumya et al., 2021). Shared-task initiatives such as LT-EDI-EACL extended hope speech detection to English, Malayalam–English, and Tamil–English, where TF-IDF features combined with MuRIL embeddings achieved F1 scores of 0.92, 0.75, and 0.57 respectively (Dave et al., 2021). Specialized datasets such as KanHope (English–Kannada) highlight persistent issues of class imbalance and preprocessing challenges involving emojis and multilingual tokens (Hande et al., 2021). Overall, methodological advances have transitioned from traditional machine learning to transformer-based architectures, where task-adaptive pre-training and multilingual contextual embeddings substantially improve performance, particularly in low-resource and morphologically rich languages (Jayanthi and Gupta, 2021; Shanmugavadivel et al., 2022). Offensive Language Identification in code-switched text presents unique challenges, as users often employ strategic language alternation to bypass keyword-based moderation. Foundational datasets such as OffMix-3L establish trilingual benchmarks for Bangla–English–Hindi, underscoring the difficulty of handling transliterated content where phonetic variation hinders detection accuracy (Goswami et al., 2023; Sazzed, 2021). Transformer-based systems such as COOLI explicitly target adversarial switching strategies, while synthetic code-switched data generation has emerged as a promising avenue for building linguistically diverse and robust training corpora (Balouchzahi et al., 2021a; Salaam et al., 2022). Recent paradigms incorporate transfer and multi-task learning, with approaches such as SetFit enabling efficient few-shot adaptation for Tamil–English detection, and multi-task frameworks demonstrating strong performance across zero-shot and fine-tuning scenarios for harmful multimodal content (Pannerselvam et al., 2024; Kumar et al., 2025).
Takeaway
LLMs have become powerful tools for multilingual hate and offensive speech detection. Specialized models such as DweshVaani demonstrate effectiveness in identifying religious hate speech in Hindi–English code-mixed contexts (Srivastava, 2025). However, safety alignment evaluations reveal persistent vulnerabilities when encountering unseen language mixture patterns, as evidenced by frameworks like Qorgau in Kazakh–Russian bilingual settings (Goloburda et al., 2025). These findings highlight the need for better-resourced datasets, robust multilingual alignment, and evaluation protocols attuned to code-switched phenomena.
Sentiment & Emotion Analysis
Sentiment and emotion analysis in CSW contexts has advanced through tasks like SemEval, with fine-tuned transformers (mBERT, XLM-R, RoBERTa) performing well on Hindi-English and Spanish-English datasets using focal loss and ensembling (Angel et al., 2020; Ma et al., 2020; Palomino and Ochoa-Luna, 2020; Sultan et al., 2020; Singh and Singh Parmar, 2020; Wu et al., 2020). Research spans Dravidian, Indonesian, Vietnamese-English, Kenyan Sheng, and trilingual settings (Chakravarthi et al., 2022; Winata et al., 2022; Van et al., 2022; Etori and Gini, 2024; Raihan et al., 2023a, 2024). Multi-label emotion detection and ABSA enhance cross-lingual performance (Wadhawan and Aggarwal, 2021; Zhang et al., 2021). Data scarcity is addressed via unsupervised self-training, curriculum learning, monolingual data, and CoSDA-ML’s synthetic augmentation (Gupta et al., 2021b; Ranjan et al., 2022; Kumar et al., 2022c; Qin et al., 2020). LLMs enable zero-shot sentiment analysis and efficient synthetic data generation (Yadav et al., 2024; Zhang et al., 2023; Zeng, 2024). Harmful content detection uses datasets for Bangla and Devanagari hate speech, with PEFT and SetFit improving results (Raihan et al., 2023b; Sidibomma et al., 2025; Pannerselvam et al., 2024).
Takeaway
Transformer-based models and LLMs, enhanced by synthetic data and few-shot learning, effectively tackle data scarcity in CSW sentiment analysis, across diverse languages (Qin et al., 2020; Pannerselvam et al., 2024). Developing robust datasets for low-resource and trilingual CSW scenarios is essential to boost generalization (Raihan et al., 2024). Lightweight adaptation techniques, like PEFT, are critical for efficient deployment in real-world multilingual applications (Sidibomma et al., 2025).
Syntactic analysis
Syntactic analysis in CSW contexts has advanced from structural modeling to linguistic theory-guided approaches, improving parsing and evaluation. SyMCoM achieved 93.4% POS tagging accuracy for English-Hindi CSW by capturing syntactic variety (Kodali et al., 2022a). Synthetic treebanks for Bengali-English dependency parsing yielded 76.24% UAS and 61.41% LAS (Ghosh et al., 2019). CoMix’s phonetic and POS-guided pretraining boosted Hinglish translation BLEU by 12.98 and NER F1 by 2.42 (Arora et al., 2023). Linguistic constraint comparisons showed a 2:1 naturalness advantage for Equivalence Constraint over heuristics ($\alpha=0.59$) (Pratapa and Choudhury, 2021). Recent work leverages LLMs for Universal Dependencies annotations, achieving 95.29 LAS after revision for Spanglish and Spanish–Guaraní CSW, bridging syntactic modeling with LLM capabilities (Kellert et al., 2025). Large-scale CSW syntax experiments reached 79.4% acceptability across 11 language pairs (Sterner and Teufel, 2025a), with pre-trained models showing 0.83 graph edit distance correlation to monolingual parses (Laureano De Leon et al., 2024). Non-English LLM prompting improved grammaticality detection by 12% to 79.6% accuracy (Behzad et al., 2024), and LLM-based GEC achieved 63.71 F0.5 on learner corpora (Potter and Yuan, 2024). Challenges remain in universalizing syntactic constraints across diverse languages (Pratapa and Choudhury, 2021).
Takeaway
Syntactic-aware models and LLMs achieve high CSW parsing performance, but universal syntactic constraints remain debated, limiting low-resource language generalization (Pratapa et al., 2018a; Behzad et al., 2024). Future work should prioritize lightweight LLM adaptations for non-Latin script and tonal languages to enhance cross-linguistic transfer (Potter and Yuan, 2024).
Intent Classification
Intent classification in code-switched conversational AI tackles irregular language alternation and non-standard orthography. Early vectorizer–classifier pipelines achieved high accuracy and robust NER (Rajeshwari and Kallimani, 2021). Random translation augmentation improved intent and slot performance on MultiATIS++ across eight languages (Krishnan et al., 2021), while XLM-R outperformed word embeddings in zero-shot Hinglish intent detection on M-CID (Arora et al., 2020). Seq2Seq models boosted slot labeling on MASSIVE, and PRO-CS prompting enhanced Hinglish accuracy (Wang et al., 2022; Bansal et al., 2022). Multilingual semantic parsing on NLmaps and ENTITYCS further improved intent and slot performance across languages (Duong et al., 2017; Chen et al., 2022), with bilingual corpora supporting joint multi-intent and entity recognition in chatbots (Aguirre et al., 2022). Recent approaches like ContrastiveMix pretraining enhanced intent accuracy across 51 languages (Lin et al., 2024). Persistent challenges include limited corpora for typologically distant languages and high computational costs of LLMs (Lin et al., 2024; Zhang et al., 2024b).
Takeaway
Code-switched intent classification has advanced with XLM-R and prompting methods like PRO-CS, but struggles with typologically diverse languages (Lin et al., 2024; Bansal et al., 2022). Innovative augmentation and lightweight transfer learning are key to overcoming data scarcity and computational barriers (Krishnan et al., 2021).
Question Answering
Code-mixed QA combines linguistic and cultural reasoning across languages. Early neural QA systems struggled with multilingual queries using synthetic monolingual data (Gupta et al., 2018). A crowd-sourced dataset improved natural code-mixing with mBERT (Chandu et al., 2018a). Multimodal visual QA with knowledge distillation enhanced efficiency in low-resource settings (Raj Khan et al., 2021). LLM-based COMMIT with instruction tuning improved Hinglish extractive QA, while non-English prompting boosted accuracy (Lee et al., 2024; Behzad et al., 2024). Curriculum learning in K-MMLU/HAE-RAE improved multiple-choice QA via code-switching pretraining (Yoo et al., 2025). Embedding transfer in ARC/TruthfulQA enhanced reasoning in Arabic, Bengali, and Vietnamese (Hong et al., 2025). MEGAVERSE showed LLMs excelling in African language QA, though test contamination affected results (Ahuja et al., 2024). Alignment methods like MAD-X addressed morphological complexity, improving Hinglish QA (Fazili and Jyothi, 2022). Still, Challenges include limited evaluation frameworks for high-resource pairs and data scarcity in non-Latin scripts (Chandu et al., 2018b; Raj Khan et al., 2021).
Takeaway
Code-mixed QA showcases robust progress with LLMs achieving up to 79.6% accuracy through informed prompting and synthetic data (Behzad et al., 2024; Gupta et al., 2018). The reliance on high-resource pairs like Hinglish reveals a gap in addressing diverse, low-resource language ecosystems, necessitating innovative data augmentation (Chandu et al., 2018a). Lightweight models are essential to democratize access for communities with natural code-switching practices (Raj Khan et al., 2021).
Natural Language Inference
Code-mixed NLI extends entailment detection across languages. The Hindi-English CALI dataset showed mBERT’s limitations due to annotation disagreements (Khanuja et al., 2020a; Huang and Yang, 2023). CoSDA-ML’s synthetic data enabled zero-shot NLI across 19 languages (Qin et al., 2020), while intermediate-task fine-tuning and meta-learning improved Hinglish NLI (Prasad et al., 2021; Kumar et al., 2022b). Prompt-based ICM boosted accuracy but requires language-specific tuning (Shankar et al., 2024). Challenges in inference and preprocessing have seen innovations that reduced API costs by 90% while maintaining performance (Nag et al., 2024). However, scaling remains difficult—mBERT and XLM-R still underperform on code-mixed tasks compared to monolingual ones (Wang et al., 2025b). CodeMixEval further reported GPT-4 Turbo’s accuracy drop in 18 low-resource languages, while MultiPragEval revealed persistent pragmatic inference issues tied to cultural context (Yang and Chai, 2025; Park et al., 2024).
Takeaway
Code-mixed NLI achieves 0.68-0.75 F1 with fine-tuned transformers and synthetic data, but cultural and pragmatic variability hinders generalization (Huang and Yang, 2023; Park et al., 2024). Future efforts should focus on diverse language-pair benchmarks and lightweight models to improve scalability (Nag et al., 2024). Efficient preprocessing and prompt engineering are critical for cost-effective, real-world deployment.
A.2 Natural Language Generation Tasks
Cross-lingual Transfer
Code-mixed cross-lingual transfer learning mitigates data scarcity by enabling models to generalize to unseen code-switched tasks. Progressive Code-Switching (PCS) incrementally augments low-resource NER data, achieving 0.82-0.87 F1 in zero-shot transfer across ten languages (Li et al., 2024). EntityCS leverages Wikidata for cross-lingual slot filling, improving SLU benchmarks by 10% F1 (Chen et al., 2022). SCOPA’s soft embedding mixing and pairwise alignment boost code-switched sentence representation by 3-5% accuracy (Lee et al., 2021). Incontext Mixing (ICM) prompts enhance intent classification by 5-8% accuracy on MultiATIS++ (Shankar et al., 2024). Test-time code-switching for aspect-based sentiment analysis yields 8-12% F1 gains on SemEval (Sheng et al., 2025). A code-switching curriculum improves intent detection by 12-15% on African languages in MASSIVE (Yoo et al., 2025). MIGRATE’s domain-specific LLM adaptation aligns embeddings, boosting zero-shot QA and NER by 15-20% F1 for low-resource languages (Hong et al., 2025). Challenges persist in handling typological diversity across language families.
Takeaway
Code-mixed transfer learning, with methods like PCS and MIGRATE, achieves 0.82-0.87 F1 by leveraging synthetic data and embedding alignment, but typological diversity remains a barrier (Li et al., 2024; Hong et al., 2025). Innovative prompting and curricula uniquely empower low-resource language tasks by simulating natural switching patterns (Shankar et al., 2024). Future efforts should focus on scalable, typologically inclusive frameworks to ensure robust generalization across diverse linguistic contexts (Yoo et al., 2025).
Text Summarization
Text summarization in code-switched contexts tackles data scarcity and linguistic complexity with specialized datasets and models. GupShup’s Hindi-English dataset (6,800+ conversations) achieved 38.2 ROUGE-L using fine-tuned mBART and multi-view seq2seq models (Mehnaz et al., 2021). CroCoSum’s 24,000 English-Chinese article-summary pairs (92% code-switched) showed that ROUGE-1 drops for cross-lingual models (Zhang and Eickhoff, 2024). CS-Sum’s Hindi-English and Spanish-English benchmark improved ROUGE-2 via alternation pattern modeling (Suresh et al., 2025). MLSUM’s multilingual corpus (five languages) gained 7% ROUGE-L in low-resource settings with synthetic data fine-tuning (Scialom et al., 2020). Contrastive learning boosted ROUGE-1 by 15% on MultiATIS++ by aligning mixed-language representations (Zhang and Eickhoff, 2024). Challenges persist in capturing semantic alignments across typologically diverse languages (Lin et al., 2024).
Takeaway
Code-switched NLG tasks everage LLMs and datasets like GupShup and CS-Sum to achieve strong results, but typological diversity and syntactic complexity at switch points remain hurdles. Techniques like PCS and contrastive learning improve low-resource tasks by aligning embeddings and replicating natural CSW patterns. Capturing cultural context and informal pragmatics is still challenging. Future efforts should prioritize adaptive models, user-centric evaluation, and typologically inclusive frameworks.
A.3 Emerging Contemporary Tasks
Code-switched dialogue generation
has emerged as a critical application area, with frameworks like MulZDG (Liu et al., 2022) enabling zero-shot dialogue generation for low-resource languages through synthetic code-mixed data augmentation. Conversational summarization for code-mixed dialogues, exemplified by the GupShup dataset (Mehnaz et al., 2021) for Hinglish conversations, represents a challenging abstractive summarization task requiring understanding of informal, mixed-language discourse patterns. Visual question answering systems have been extended to handle code-mixed queries through knowledge distillation approaches that transfer capabilities from large multilingual models to efficient specialized architectures (Raj Khan et al., 2021). Code-mixed text quality evaluation has emerged as a specialized task, with approaches combining linguistic metrics and multilingual language model embeddings to assess synthetic data adequacy and fluency (Kodali et al., 2022b). Homophonic pun generation in code-mixed contexts demonstrates the creative potential of LLMs in cross-linguistic wordplay and humor generation (Sarrof, 2025).
A.4 Underexplored Frontiers Tasks
Although notable progress has been made in core code-switching tasks, several frontier areas remain underexplored. Reasoning tasks such as mathematical, and causal reasoning struggle with cross-language entailment and logical complexity (Raihan et al., 2023a; Mohamed et al., 2025). Higher-order cognitive tasks like metaphor understanding and analogical reasoning show limited performance and lack cultural grounding (Kodali et al., 2025; Mehnaz et al., 2021). Coding-related tasks, including code generation and explanation from mixed prompts, achieve moderate functional accuracy but are far from mature (Yang and Chai, 2025; Khatri et al., 2023). While controllable language steering has shown early success (e.g., 85% switch-point control) (Mondal et al., 2022), adaptive and pragmatic CSW generation remains largely theoretical (Kuwanto et al., 2024).
Takeaway
Several challenges remain: (i) data sparsity and imbalance, especially for underrepresented languages; (ii) orthographic variability and noisy social media text that reduce generalizability; (iii) limited scalability of transfer learning methods for distant language pairs; (iv) insufficient exploration of shared versus task-specific representations in joint modeling; and (v) lack of robust, community-wide benchmarks for fair comparison across diverse CSW scenarios.
A.5 Fine-tuning Approaches
Instruction Tuning
Instruction tuning in multilingual (CSW) settings enhances LLMs ability to follow instructions across languages while aligning with human preferences, despite challenges like Script variabilty and cultural nuances. COMMIT adapts English-centric LLMs via code-mixed instruction tuning for low-resource QA, achieving up to 32× gains in exact match on Hinglish tasks (Lee et al., 2024). CSCL progressively introduces CSW during instruction tuning, improving cross-lingual transfer by 5–8% across 11 language pairs (Yoo et al., 2025). SPHINX achieves sample-efficient fine-tuning via translated instructions, boosting zero-shot QA performance by 10–15% in African languages (Ahuja et al., 2025). PLUG uses pivot-language code-switching to enhance instruction-following, yielding 7–12% gains in multilingual response generation (Zhang et al., 2024c). Preference-aligned methods, including Predicting Preferences in Generated CSW Text and Multilingual Blending, improve naturalness and ethical adherence by 15–20% across low-resource scenarios (Gupta et al., 2025; Song et al., 2025). These works highlight the impact of curriculum-based and preference-optimized tuning, while underscoring the need for culturally diverse datasets.
Takeaways:
Curriculum and preference-guided fine-tuning enables LLMs to better handle multilingual CSW tasks with improved instruction adherence and cross-lingual generalization. Future research should emphasize robustness to spontaneous CSW, cross-domain generalization, and contextually appropriate outputs.
Parameter-efficient fine-tuning
PEFT methods like LoRA, QLoRA, and soft prompt tuning efficiently adapts LLMs for CSW tasks. PEFT methods like LoRA on Llama-3.1-8B and Nemo-Instruct achieves up to 90.75% accuracy for Hindi and Nepali hate speech (Sidibomma et al., 2025), QLoRA on Gemma-2 reaches 76.19% for Hinglish religious hate speech (Srivastava, 2025), and soft prompt tuning reduces MER by 10%+ in Mandarin-English speech recognition (Liu et al., 2025). LoRA improves Hinglish NER by 5–10% despite transliteration challenges (Shirke et al., 2025). Adapters and quantization-aware PEFT lower resource demands for safety evaluation and constrained generation but require careful hyperparameter tuning (Goloburda et al., 2025; Kuwanto et al., 2024). Overall, PEFT provides scalable, resource-efficient adaptation for code-switched LLMs (Weyssow et al., 2024, 2024).
Takeaways:
PEFT techniques like LoRA, QLoRA, and soft prompt tuning efficiently adapt LLMs to code-switched and low-resource multilingual settings with minimal parameters and hardware demands. They consistently improve performance in tasks like hate speech detection, NER, and speech recognition, while reducing computational cost.
Reinforcement Learning for CSW Adaptation
To improve LLMs’ code-mixing understanding, reinforcement learning from AI feedback (RLAIF) has been proposed as a cost-efficient alternative to human annotation, effectively enhancing code-mixed translation (Zhang et al., 2023). CHAI applies RLAIF on Llama-3.1-8B-Instruct for English-Hinglish translation, generating preference pairs via GPT-4o on 15,190 MixMT and ALL-CS instances, and uses PPO for 5 epochs to maximize reward while controlling KL divergence, achieving 40–46% higher human-evaluated win rates, 24–81% gains in chrF and COMET translation metrics, and 14.12% accuracy improvement in Hinglish sentiment analysis (Zhang et al., 2025c). Similarly, (Zhang et al., 2025c) enhances multilingual LLMs’ handling of code-mixing using RLAIF and code-mixed machine translation, reducing reliance on human annotations by leveraging LLMs for preference labeling. (Heredia et al., 2025b) employs RL-based policy optimization with back-translation of naturally occurring CS sentences to create synthetic data, improving generation naturalness and fluency by optimizing for acceptability scores (Heredia et al., 2025b).These methods enhance CSW fluency and alignment, though typological diversity and computational cost remain challenges (Zhang et al., 2025c).
Takeaway
RLAIF and RL-based methods for synthetic data generation represent promising but underexplored approaches to enhance LLMs code-mixed understanding while reducing reliance on human annotations.
Supplementary Material:
This section provides additional resources to support our main findings, including extended tables, illustrative examples of model hallucinations, and dataset analyses for code-mixed NLP research.
Figure 3: Failure cases when we prompt ChatGPT in Odia-Romanized Hindi code-mixed pair.
Figure 4: Failure cases when we prompt GLM-4.6 in Bangla-English code-mixed pair.
The image 3 shows a ChatGPT hallucination where a query for vacation countries results in a list of Indian states and union territories. The response also exhibits incorrect grammar, demonstrating the model’s limitations in handling nuanced, code-mixed queries.
The image 4 shows a failure case from Perplexity AI. When prompted in a code-mixed query for Goan temples specifically destroyed by the Portuguese, the model disregards this critical constraint and instead provides information on existing ancient temples.
Figure 5: Failure cases when we prompt Perplexity in Konkani-English code-mixed pair.
The image 5 shows an illustration of hallucination in the open-source model GLM-4.6. When prompted to perform morphophonemic (sandhi) analysis of the Bengali word Ohoroho (“always”) in Bengali-Romanized Hindi script, the model misinterprets the query and generates an incorrect response.
Figure 6: Analysis of Code-Switching Dataset Sizes.
Figure 6 presents a dual analysis of dataset sizes. The bar chart on the left compares the number of instances in prominent datasets, showing COMI-LINGUA (125,615) and MEGAVERSE (100,000) as the largest. The pie chart on the right categorizes the overall distribution, revealing that ’Large’ datasets (50-100k instances) are the most common category, comprising 40% of the analyzed collections.
Figure 7: Top 20 Language Pairs in Code-Switching Research
Figure 8: Distribution of NLP Tasks and Categories in Code-Switching Research.
Figure 7 chart displays the prevalence of different language pairs in code-switching studies, measured by the number of available datasets and benchmarks. The data clearly indicates a strong dominance of Hindi-English and Spanish-English combinations, highlighting a significant research focus on these pairs compared to others.
Figure 8 illustrates the primary focus areas within code-switching NLP research. The left pie chart details the distribution of specific tasks, with SA (26.2%) and MT (18.5%) being the most studied. The right pie chart groups these into broader categories, where ’Other’ (55.9%) and ’Understanding’ (31.1%) tasks represent the vast majority of research efforts.
Figure 9 shows the distribution of language pairs in 202 code-mixing related datasets and benchmarks papers, highlighting the prevalence of specific language combinations in code-switching NLP research.
Figure 9: Language pair distribution across 202 code-mixing related datasets and benchmarks papers.
Category Tasks/Components Datasets & Benchmarks Model Training & Adaptation Evaluation & Methodology Key Insights
NLU Tasks: Foundational LID (MasakhaNER, ECT); POS tagging (XLM-R, IndicBERT) (Absar, 2025) ; NER (MELM, GLUECoS) (Zhou et al., 2022; Khanuja et al., 2020b); Syntactic parsing (CoStALL LLM) LinCE (Aguilar et al., 2020), GLUECoS (Khanuja et al., 2020b), MultiCoNER (Malmasi et al., 2022b), PACMAN (50K+ samples) (Chatterjee et al., 2022), WILDRE (Kumar et al., 2022a) XLM-RoBERTa (F1: 0.88) (Kochar et al., 2024) ; MuRIL (5-8% > mBERT) (Goswami et al., 2023); IndicBERT (F1: 0.82) (Tatariya et al., 2023); Traditional: CRF, SVM Benchmarks: GLUECoS (6 tasks) (Khanuja et al., 2020b), LinCE (Aguilar et al., 2020); Metrics: F1, accuracy; Cross-lingual transfer XLM-R on MultiCoNER (F1: 0.88) (Malmasi et al., 2022b); MuRIL (Goswami et al., 2023) best for Indic; handles script mixing natively
NLU Tasks: QA & Understanding Question Answering (COMMIT) (Lee et al., 2024); Information Retrieval; Reading Comprehension; Semantic Understanding COMMIT-based datasets (Lee et al., 2024); MLQA (multilingual QA) (Lewis et al., 2020); Low-resource QA benchmarks COMMIT (LLaMA-7B) (Lee et al., 2024): 32x improvement on MLQA (Lewis et al., 2020); Code-mixed instruction tuning; Template preservation with content mixing Metrics: Exact Match, F1; Cross-lingual transfer; Few-shot evaluation COMMIT (Lee et al., 2024) 32x improvement on MLQA (Lewis et al., 2020); demonstrates power of code-mixed instruction tuning
NLG Tasks MT (CoMix) (Arora et al., 2023); Summarization (CroCoSum) (Zhang and Eickhoff, 2024); Dialogue Generation (X-RiSAWOZ: 18K+ utterances, 12 domains) (Moradshahi et al., 2023); Creative text (puns); CS prompting CroCoSum (24K En articles, 18K Chinese summaries) (Zhang and Eickhoff, 2024); X-RiSAWOZ (En, Fr, Hi, Ko, En-Hi) (Moradshahi et al., 2023); IndioTrans; HinglishEval (Srivastava and Singh, 2022b) CoMix (Arora et al., 2023): BLEU 12.98, 10x smaller than mT5; GPT-4 strong zero-shot generation; Flan-T5-XXL (Yong et al., 2023) limited CS capability Metrics: BLEU, ROUGE, CS-MI; Task-specific: DST, RG; Quality estimation (HinglishEval) (Srivastava and Singh, 2022b); Integration: transliteration, Indic-TTS CoMix (Arora et al., 2023) BLEU: 12.98 with 10x efficiency; GPT-4 excels at creative generation (puns, translation)
NLU Tasks: Classification Sentiment (DravidianCodeMix: 60K+ comments) (Chakravarthi et al., 2022); Offensive detection (OffMix-3L) (Goswami et al., 2023); Hate speech; Intent classification DravidianCodeMix (Tamil-En, Kannada-En, Malayalam-En) (Chakravarthi et al., 2022); OffMix-3L (Goswami et al., 2023); SentMix-3L (Raihan et al., 2023a); MEGAVERSE (83 languages) (Ahuja et al., 2024) Sentence-BERT SetFit (F1: 0.72, minimal data) (Pannerselvam et al., 2024); mBERT (F1: 0.85+ baseline); GPT-4 zero-shot strong MEGAVERSE (Ahuja et al., 2024) holistic evaluation; Metrics: perplexity, grammaticality; Social media analysis workflows SetFit (Pannerselvam et al., 2024) F1: 0.72 with minimal data for Dravidian (Chakravarthi et al., 2022); GPT-4 strong on SentMix/OffMix-3L (Raihan et al., 2023a; Goswami et al., 2023)
Specialized Models & Techniques Pre-training: HingBERT (L3Cube-HingCorpus) (Nayak and Joshi, 2022); Instruction tuning; Domain adaptation; POS-guided attention Synthetic: SynCS, LLM-generated; Low-resource: WILDRE (Magahi-En, Maithili-En) (Rani et al., 2024b); Language-specific corpora HingBERT (Hinglish specialist, Roman script tokenizer) (Nayak and Joshi, 2022); CoMix (POS-guided + phonetic signals) (Arora et al., 2023); Flan-T5-XXL (baseline, fails at CS) (Yong et al., 2023) Pre-training metrics; Multi-task adaptation; Phonetic feature integration; Zero-shot vs fine-tuned comparison HingBERT outperforms mBERT on Hinglish; CoMix 10x efficiency; specialized tokenization critical
Benchmarks & Evaluation Foundational: LID, POS, NER; Understanding: sentiment, QA; Generation: MT, summarization; Novel: code generation (CodeMixBench) (Sheokand et al., 2025) COMI-LINGUA (125K instances, Roman+Devanagari) (Sheth et al., 2025); SwitchLingua (next-gen diversity) (Xie et al., 2025); HinglishEval (quality estimation) (Srivastava and Singh, 2022b); CodeMixBench (programming from CS) (Sheokand et al., 2025) Evaluation across XLM-R, MuRIL, IndicBERT, GPT-4, mBERT; Zero-shot vs fine-tuned performance Traditional: F1, BLEU; Task-specific: Prothaba; Intrinsic: perplexity; Extrinsic: GAME; Quality estimation for synthetic CS MEGAVERSE (83 langs) (Ahuja et al., 2024); COMI-LINGUA (Hinglish) (Sheth et al., 2025); CodeMixBench (Sheokand et al., 2025) introduces novel programming task
Language Coverage & Gaps High-resource: Hindi-En, Spanish-En; Dravidian: Tamil-En, Kannada-En, Malayalam-En; Low-resource: Magahi-En, Maithili-En; East Asian: Chinese-En SwitchLingua (Xie et al., 2025) (multi-ethnic); X-RiSAWOZ (5 languages) (Moradshahi et al., 2023); MuRIL corpus (17M Indian) (Goswami et al., 2023); MultiCoNER (11 languages) (Malmasi et al., 2022b) Universal: XLM-R (100 langs), mBERT (104 langs); Regional: MuRIL (17 Indic), IndicBERT (12 Indic) (Goswami et al., 2023); GPT-4 (inconsistent on low-resource) Cross-lingual transfer; Probing tasks; Zero-shot assessment; Task-oriented dialogue evaluation (X-RiSAWOZ) (Moradshahi et al., 2023) Low-resource pairs underserved; script mixing only by specialists; unseen CS patterns challenge all models
Table 1: Comprehensive landscape of code-switching NLP: Tasks, datasets, models, and evaluation across all categories. Takeaway (1) Specialized models (MuRIL, IndicBERT) outperform general-purpose on specific language families; (2) Instruction tuning yields dramatic gains (COMMIT: 32x); (3) Efficiency matters (CoMix: 10x smaller, better performance); (4) Few-shot learning viable (SetFit: F1 0.72 minimal data); (5) Major gaps: low-resource languages, unseen CS patterns, generation consistency.
Task Dataset Languages Domain Key Characteristics
NER SemEval-2022 Task 11 (Malmasi et al., 2022a) Multilingual Short queries 6 entity types (Person, Location, Group, Corp, Product, Creative Work)
Kannada-English NER (S and Shrivastava, 2022) Kannada-English Social media Low-resource Dravidian language; user-generated content
TB-OLID (Raihan et al., 2023b) Bangla-English Social media 5,000 Facebook comments; transliterated text; hierarchical annotation
[…]
CoMeT Corpus (Gautam et al., 2021b) Multiple pairs General Synthetic generation using parallel monolingual sentences
AfroCS-xs (Olaleye et al., 2025) 4 African + EN Agriculture Human-validated synthetic data for low-resource African languages
CoVoSwitch (Kang, 2024) Multiple pairs Synthetic/Speech Prosody-aware synthetic data generation for MT based on intonation units
Dialogue GupShup (Mehnaz et al., 2021) Hindi-English Entertainment First code-mixed dialogue summarization; movie-based conversations
Multilingual WOZ 2.0 (Hung et al., 2022) Multilingual Restaurant Cross-lingual dialogue state tracking benchmark
THAR (Srivastava, 2025) Hindi-English Social media 11,549 YouTube comments; religious hate speech detection
Emotion/ Sentiment EmoMix-3L (Raihan et al., 2024) Bangla-EN-Hindi Social media 1,071 instances; 5 emotion classes (Happy, Surprise, Neutral, Sad, Angry)
SentMix-3L (Raihan et al., 2023a) Bangla-EN-Hindi Social media 1,007 instances; 3-class sentiment; controlled trilingual collection
OffMix-3L (Goswami et al., 2023) Bangla-EN-Hindi Social media 1,001 instances; offensive language identification; first trilingual dataset
ASR/Speech ASCEND (Lovenia et al., 2022) Mandarin-English Conversational Hong Kong speakers; location-specific variation documented
SEAME (Lyu et al., 2010) Mandarin-EN-Hokkien Conversational Singapore speakers; includes three-language mixing
English-isiZulu CS (Biswas et al., 2020) English-isiZulu Conversational Semi-supervised acoustic and language modeling; low-resource African
Table 2: Top specialized datasets by task category.
Model / Method Type Target Scenario Key Innovation & Results
HingBERT (Nayak and Joshi, 2022) Pre-trained encoder Hindi-English (Hinglish) Pre-trained on L3Cube-HingCorpus; outperforms mBERT on Hinglish sentiment, NER; specialized tokenizer for Roman script Hindi
COMMIT (LLaMA-7B based) (Lee et al., 2024) Instruction-tuned LLM Low-resource QA Code-mixed instruction tuning; 32x improvement on MLQA; specialized code-mixing preserves templates while mixing content
CoMix (BERT/BART variants) (Arora et al., 2023) Domain-adapted transformer Hinglish translation, NER POS-guided attention; phonetic signals; SOTA on LINCE (BLEU: 12.98); 10x smaller than mT5 with better performance
Flan-T5-XXL (Yong et al., 2023) Multilingual LLMs Zero-shot generation Limited code-mixing capability; fails at code-mixing; useful baseline but needs fine-tuning for most tasks
Sentence-BERT (SetFit) (Pannerselvam et al., 2024) Few-shot classifier Low-resource offensive detection Achieves F1: 0.72 with minimal labeled data; efficient for Tamil, Telugu offensive content; practical for data-scarce scenarios
Table 3: Specialized models and adaptation techniques for code-switching.
Model Methodology (brief) Strengths Weaknesses
XLM-RoBERTa (Kochar et al., 2024) Multilingual masked LM trained on 2.5TB CommonCrawl; 100 languages; RoBERTa architecture + SOTA on MultiCoNER (F1: 0.88), OffMix-3L; excellent zero-shot transfer; robust cross-lingual representations Struggles with unseen code-switching patterns; requires fine-tuning for best results; computationally expensive
MuRIL (Goswami et al., 2023) BERT pre-trained on 17 Indian languages + transliterated text ; 17M Indian corpus + Best for Indic tasks; handles script mixing; 5-8% better than mBERT on Hindi-English Limited to Indian subcontinent; less effective for other language families; smaller coverage
mBERT (Goswami et al., 2023) Multilingual BERT on 104 Wikipedia dumps; shared vocabulary + Strong baseline (F1: 0.85+); widely adopted; stable performance across tasks Curse of multilinguality; undertrained on low-resource languages; outperformed by specialized models
GPT-4 (Ahuja et al., 2024) Decoder-only transformer; web-scale training; RLHF alignment + Strong zero-shot on SentMix-3L, OffMix-3L; excellent generation (puns, translation); few-shot learning Closed-source; expensive API; inconsistent on low-resource pairs; unpredictable behavior
IndicBERT (Tatariya et al., 2023) BERT on 12 Indian languages; 9GB Indic corpus; language-specific tokenization + Best for Indian monolingual + code-mixed tasks; F1: 0.82 on DravidianCodeMix; efficient Limited to 12 languages; requires language-specific tuning; less generalizable than XLM-R
Table 4: Top 5 models for code-switching NLP with methodology and performance characteristics.
Benchmark Description Language Pairs Key Focus
LinCE (Aguilar et al., 2020) A centralized benchmark that aggregated several existing datasets (e.g., for LID, POS, NER) to standardize evaluation of foundational linguistic tasks. Spanish-En, Nepali-En, Hindi-En, etc. Foundational NLU Tasks
GLUECoS (Khanuja et al., 2020b) The first multi-task benchmark suite for code-switching, modeled after the influential GLUE benchmark for English, covering 6 diverse tasks. Hindi-En, Spanish-En Multi-task NLU Evaluation
DravidianCodeMix (Chakravarthi et al., 2022) Large-scale dataset with over 60,000 YouTube comments for sentiment analysis and offensive language identification in Dravidian languages. Tamil-En, Kannada-En, Malayalam-En Sentiment & Offensive Language
MultiCoNER (Malmasi et al., 2022b) A large-scale multilingual benchmark for "complex" NER, featuring nested entities and code-mixing across 11 languages. 11 languages Complex NER
PACMAN (Chatterjee et al., 2022) A synthetically generated code-mixed POS-tagged dataset with over 50K samples, outperforming existing benchmarks in code-mixed POS tagging. Hindi-En POS Tagging
WILDRE Shared Tasks (Kumar et al., 2022a) A series of shared tasks focused on promoting research for less-resourced Indo-Aryan languages on tasks like sentiment analysis. Magahi-En, Maithili-En, etc. Low-Resource Languages
CroCoSum (Zhang and Eickhoff, 2024) A dataset for cross-lingual code-switched summarization of technology news, with over 24,000 English source articles and 18,000 human-written Chinese summaries. English–Chinese Cross-lingual Summarization
X-RiSAWOZ (Moradshahi et al., 2023) A multi-domain, large-scale, high-quality task-oriented dialogue benchmark created by translating the Chinese RiSAWOZ dataset to English, French, Hindi, Korean, and a code-mixed English-Hindi language. Contains over 18,000 utterances per language across 12 domains. En, Fr, Hi, Ko, En-Hi (code-mixed) End-to-end Task-oriented Dialogue (DST, ACD, DAG, RG)
SwitchLingua (Xie et al., 2025) A recent, large-scale multilingual and multi-ethnic dataset designed to be the next generation of code-switching benchmarks. Multiple Diversity & Inclusivity
MEGAVERSE (Ahuja et al., 2024) A massive-scale, multi-modal benchmark covering 83 languages, designed for a holistic and challenging evaluation of modern LLMs. 83 languages Broad LLM Capabilities
HinglishEval (Srivastava and Singh, 2022b) A generation challenge focused on the difficult task of Quality Estimation for synthetically generated code-mixed text. Hindi-En Generative Evaluation
CodeMixBench (Sheokand et al., 2025) The first benchmark for the novel and non-traditional task of generating programming code from code-mixed natural language prompts. Hindi-En, Spanish-En Non-traditional Tasks
COMI-LINGUA (Sheth et al., 2025) Expert-annotated 1,25,615 Hinglish code-mixed instances across five core NLP tasks, covering both Roman and Devanagari scripts. Hi-En Multi-task: LID, MLI, POS, NER, MT
Table 5: An expanded overview of major benchmarks for standardized evaluation of code-switching NLP models.
Dataset Description Language Pairs Tasks
Hinglish Blog Corpus (Chandu et al., 2018b) A corpus from eight Hinglish blogging websites covering diverse topics, containing 59,189 unique sentences. Hindi-En POS Tagging, Language Modeling
CS-NLI (Chakravarthy et al., 2020) A dataset for Hinglish conversational entailment derived from Bollywood movie scripts, containing 2,240 premise-hypothesis pairs. Hindi-En Natural Language Inference (NLI)
Bollywood NLI (Khanuja et al., 2020a) A conversational NLI dataset from 18 Bollywood movie scripts, with 400 premises and 2,240 hypotheses. Hindi-En Natural Language Inference (NLI)
DravidianCodeMix (Chakravarthi et al., 2022) A manually annotated dataset for sentiment and offensive language detection in Dravidian languages. Contains over 60,000 sentences. Tamil-En, Mal.-En, Kan.-En Sentiment, Offensive ID
HinGE (Srivastava and Singh, 2021) A high-quality, manually annotated Hindi-English dataset for NLG tasks, with 3,872 sentences annotated by five annotators. Hindi-En Natural Language Generation (NLG)
Bengali Abusive Comments (Sazzed, 2021) A balanced corpus of 3,000 YouTube comments (1,500 abusive, 1,500 non-abusive) in transliterated Bengali-English. Bengali-En Abusive Language Detection
MultiCoNER (Malmasi et al., 2022b) A large-scale multilingual benchmark for "complex" NER, featuring nested entities and code-mixing across 11 languages. 11 languages Complex Named Entity Recognition
ASCEND (Lovenia et al., 2022) A large-scale dataset (10.3 hours) of spontaneous, multi-turn spoken conversational code-switching, capturing natural dialogue. Mandarin-En Dialogue, ASR
GupShup (Mehnaz et al., 2021) The first large dataset for abstractive summarization of informal, code-switched conversations. Contains 6,800 conversations. Hindi-En Dialogue Summarization
TweetTaglish (Imperial and Ong, 2022) A large dataset of 78,689 Tagalog-English code-switched tweets, expanding resources for Southeast Asian languages. Tagalog-En Language ID, etc.
My Boli (Joshi et al., 2023) A comprehensive suite of corpora and pre-trained models for the under-resourced Marathi-English language pair. Marathi-En General NLU
OffMix-3L / EmoMix-3L (Goswami et al., 2023; Raihan et al., 2023b) The first datasets for affective tasks involving a mix of three languages simultaneously, created from social media. Bangla-En-Hi Offensive ID, Emotion
EkoHate (Ilevbare et al., 2024) A dataset of 3,398 tweets for hate speech detection in Nigerian English and Pidgin political discussions. Nigerian En-Pidgin Hate Speech Detection
MaCmS (Rani et al., 2024a) The first Magahi-Hindi-English code-mixed dataset for sentiment analysis, focusing on a less-resourced minority language. Magahi-Hindi-En Sentiment Analysis
Prabhupadavani (Kumar et al., 2024) A massive and unique code-mixed speech translation dataset covering 25 languages, derived from religious discourses. 25 Indic Lang.-En Speech Translation
KRCS (Borisov et al., 2025) The first Kazakh-Russian code-switching parallel corpus for machine translation, containing 618 parallel sentences. Kazakh-Russian Machine Translation
BarNER (Peng et al., 2024) The first dialectal NER dataset for German, containing 161,000 tokens annotated from Bavarian Wikipedia and tweets. Bavarian-German Named Entity Recognition (NER)
BanglishRev (Shamael et al., 2024) A large-scale (23,546 reviews) Bangla-English code-mixed dataset of e-commerce product reviews. Bengali-En Sentiment Analysis
BnSentMix (Alam et al., 2025) A large (>20,000 sentences) code-mixed Bengali dataset from diverse sources (Facebook, YouTube, e-commerce). Bengali-En Sentiment Analysis
MMS-5 (Arul et al., 2025) A multi-modal, multi-scenario hate speech dataset for Dravidian languages, including code-mixed text and images. Tamil-En, Kannada-En Multimodal Hate Speech
Qor’gau (Goloburda et al., 2025) The first comprehensive safety evaluation dataset for bilingual Kazakh-Russian contexts, with 500 code-switched examples. Kazakh-Russian LLM Safety Evaluation
AfroCS-xs (Olaleye et al., 2025) A compact, human-validated synthetic dataset for four African languages, demonstrating the power of high-quality data over quantity. 4 African Lang.-En Machine Translation
DweshVaani (Srivastava, 2025) An LLM fine-tuned on Gemma-2 for detecting religious hate speech in code-mixed Hindi-English. Hindi-En Religious Hate Speech Detection
Cline (Kodali et al., 2025) The largest dataset of human acceptability judgments for Hinglish code-mixed text, showing low correlation with traditional metrics. Hindi-En Acceptability Judgments
Hindi-Marathi CS Corpus (Hemant and Narvekar, 2025) A 450-hour annotated code-switched speech dataset covering various switching patterns for ASR and LID. Hindi-Marathi ASR, Language ID
Word-Level Hate Speech (Niederreiter and Gromann, 2025) A multilingual dataset for word-level hate speech detection, exploring domain transfer across code-mixed language pairs. Hi-En, De-En, Es-En Word-Level Hate Speech
Table 6: An expanded overview of major datasets for code-switching NLP, highlighting language pairs, tasks, and scale.